# Disentangling Representations in RNNs through Multi-task Learning

### v1

@@ Appendix B Theoretical Derivations
Figure S2: Bayesian graphical model framework representing our theoretical framework for multi-task classification. The agent with latent state $\mathbf{Z}(t)$ estimates the ground truth decision output $\mathbf{y}(\mathbf{x^{*}})\in\{0,1\}^{N_{\text{task}}}$ from noisy observations $\mathbf{X}(t)$ transformed by injective observation map $f$. We prove that latent state $\mathbf{Z}(t)$ must encode an optimal, disentangled estimate of the de-noised environment state $\mathbf{x^{*}}$ when the decision boundary normal vectors $\{\mathbf{c}_{i}\}_{i=1}^{N_{\text{task}}}$ span $\mathbb{R}^{D}$.
@@
  • $N_{task}$ is the number of classification tasks,
  • $\{(\mathbf{c}_{i},b_{i})\}_{i=1}^{N_{task}}$ are the classification boundary normal vectors and offsets respectively, with $\mathbf{c}_{i}\in\mathbb{R}^{D}$ and $b_{i}\in\mathbb{R}$. We assume each $\|\mathbf{c}_{i}\|=1$.
  • $(\mathbf{C},\mathbf{b})$ are a matrix and vector representing each of the $N_{task}$ classification tasks where $\mathbf{C}\in\mathbb{R}^{N_{task}\times D}$
  • $\mathbf{y}(\mathbf{x}^{*})\in\{-1,+1\}^{N_{task}}:$ Ground truth classification outputs, where each ground truth classification $y_{i}(\mathbf{x}^{*})$ is given by
    $y_{i}(\mathbf{x})=\begin{cases}+1&\text{ if }\mathbf{c}_{i}^{\top}\mathbf{x}>b_{i}\\ -1&\text{ otherwise }\end{cases}$ (8)
@@
  • $g:$ Map from latent state $\mathbf{Z}(t)$ to multi-task classification estimates $\hat{\mathbf{Y}}(t)$. For our RNN experiments, readout map $g=\tanh$, for instance.
  • $\hat{\mathbf{Y}}(t):=g(\mathbf{Z}(t))\in[0,1]^{N_{task}}:$ Output vector of the multi-task classification model at time $t$, where each $\hat{Y}_{i}(t)$ is a Bernoulli random variable estimating the conditional probability $\Pr\{y_{i}(\mathbf{x}^{*})=+1\}$ given the noisy observations (via latent variable $\mathbf{Z}(t)$ – see Equation 9).
@@
We prove in Theorem B.6 (“Optimal Representation Theorem”) that any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ described above will represent an optimal estimate of $\mathbf{x}^{*}$ in latent state $\mathbf{Z}(t)$. We begin by proving results on optimal estimators in Sections B.1B.2 with identity observation map $f$, developing the linear case of the optimal representation theorem (Theorem B.5) showing that the latent state $\mathbf{Z}(t)$ must encode an estimate of $\mathbf{x}^{*}$ (visualized in Figure S3). We generalize this result any injective observation map $f$ in Section B.3 and derive closed-form solutions for extracting the estimate of $\mathbf{x}^{*}$ from $\mathbf{Z}(t)$. We derive approximation results for $g=\tanh$ in Corollary B.7 that show the representation of $\mathbf{x}^{*}$ in $\mathbf{Z}(t)$ will be disentangled (i.e., linear-affine decodable) if $g=\tanh$.
@@ B.2 Trilateration via Multiple Decision Boundaries
Let $\hat{\mathbf{Y}}(t)$ denote the vector of classification estimates $\hat{Y}(t)$ from Equation 12. We now have the tools to prove our final result via trilateration. Much like distance information from cell towers can be used to trilaterate [^2] Trilateration differs from triangulation, and it is more frequently used in practice. Triangulation is when one has angle information w.r.t. the cell towers. Usually, this is not available – so one trilaterates their position [56]. This more closely matches our setting, where we just have distances information w.r.t. the decision boundaries and must determine the position. one’s position, we will leverage Lemma B.3 and use distances from decision boundaries $\{(\mathbf{c}_{i},b_{i})\}_{i\in[N_{task}]}$ to constrain the positions.
@@
Theorem B.5 boils down to the observation that the confidence associated with each $\hat{Y}_{i}$ in $\hat{\mathbf{Y}}(t)$ are measures of distance between an implied estimate of $\mathbf{x}^{*}$ (denoted $\mu(t)$) and classification boundary $i$ (denoted $(\mathbf{c}_{i},b)$). $\hat{\mathbf{Y}}$ specifies the position of $\hat{\mathbf{X}}=\mu$ via “coordinates” defined by decision boundary normal vectors $\mathbf{c}_{1},\dots,\mathbf{c}_{N_{task}}$.
@@ B.3 Optimal Representation Theorem (General Case)
Let $\mathbf{x}^{*}\in\mathbb{R}^{D}$ be a latent disentangled representation for linear binary classification task $\mathbf{y}(\mathbf{x}^{*})\in\{0,1\}^{N_{task}}$ and $\mathbf{X}(t)=f(\mathbf{x}^{*}+\sigma\mathcal{N}(\mathbf{0},\mathbf{I}_{D}))$ be noisy observations transformed by an injective observation map $f$.
If $\mathbf{C}\in{N_{task}\times D}$ is a full-rank matrix representing the decision boundary normal vectors in $\mathbb{R}^{D}$ and $N_{task}\geq D$, then any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode an optimal representation of the latent variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$. Furthermore, the mutual information between $\mathbf{x}^{*}$ and $\mathbf{Z}(t)$ is at least as great as the entropy of the theoretically optimal estimator $\mu(t)$, ensuring that all the information of $\mu(t)$ is contained in $\mathbf{Z}(t)$. Consequently, $\mu(t)$ – the optimal estimate of $\mathbf{x}^{*}$ based on $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$ – can be written as a deterministic function (Equation 17) of latent state $\mathbf{Z}(t)$.
@@
Motivated by the similarity between $\Phi(z)$ and $tanh(z)$, we show that the two can be approximately canceled in Equation 17, implying that $\mu(t)$ can be reconstructed with high accuracy with a linear-affine transformation (e.g., linear decoding) when $g=\tanh$. This implies that $\mathbf{Z}(t)$ contains a linear disentangled representation of $\mu(t)$ [30].
@@
Page [58] proposed a simple approximation of $\Phi$ via $\tanh$. Eliminating higher order terms, their approximation is $\Phi(x)\approx\frac{1}{2}\tanh(\sqrt{\frac{2}{\pi}}x)+\frac{1}{2}$. We found the following first-order approximation yielded a superior mean squared error:
@@
Therefore, we have shown that $\mu(t)$ can be expressed as a linear transformation of $\mathbf{Z}(t)$ when the readout function $g$ is $\tanh$. This confirms the corollary.
B.4 Discussion

### v2

@@ Appendix B Theoretical Derivations
Figure S5: Bayesian graphical model framework representing our theoretical framework for multi-task classification. The agent with latent state $\mathbf{Z}(t)$ estimates the ground truth decision output $\mathbf{y}(\mathbf{x^{*}})\in\{0,1\}^{N_{\text{task}}}$ from noisy observations $\mathbf{X}(t)$ transformed by injective observation map $f$. We prove that latent state $\mathbf{Z}(t)$ must encode an optimal, linearly decodable estimate of the de-noised environment state $\mathbf{x^{*}}$ when the decision boundary normal vectors $\{\mathbf{c}_{i}\}_{i=1}^{N_{\text{task}}}$ span $\mathbb{R}^{D}$.
@@
  • $N_{\textit{task}}$ is the number of classification tasks,
  • $\{(\mathbf{c}_{i},b_{i})\}_{i=1}^{N_{\textit{task}}}$ are the classification boundary normal vectors and offsets respectively, with $\mathbf{c}_{i}\in\mathbb{R}^{D}$ and $b_{i}\in\mathbb{R}$. We assume each $\|\mathbf{c}_{i}\|=1$.
  • $(\mathbf{C},\mathbf{b})$ are a matrix and vector representing each of the $N_{\textit{task}}$ classification tasks where $\mathbf{C}\in\mathbb{R}^{N_{\textit{task}}\times D}$
  • $\mathbf{y}(\mathbf{x}^{*})\in\{0,1\}^{N_{\textit{task}}}:$ Ground truth classification outputs, where each ground truth classification $y_{i}(\mathbf{x}^{*})$ is given by
    $y_{i}(\mathbf{x})=\begin{cases}1&\text{ if }\mathbf{c}_{i}^{\top}\mathbf{x}>b_{i}\\ 0&\text{ otherwise }\end{cases}$ (11)
@@
  • $g:$ Map from latent state $\mathbf{Z}(t)$ to multi-task classification estimates $\hat{\mathbf{Y}}(t)$. For most of our experiments, readout map $g=\text{sigmoid}$, for instance.
  • $\hat{\mathbf{Y}}(t):=g(\mathbf{Z}(t))\in[0,1]^{N_{\textit{task}}}:$ Output vector of the multi-task classification model at time $t$, where each $\hat{Y}_{i}(t)$ is a Bernoulli random variable estimating the conditional probability $\Pr\{y_{i}(\mathbf{x}^{*})=1\}$ given the noisy observations (via latent variable $\mathbf{Z}(t)$ – see Equation 12).
@@
We prove in Theorem B.6 (“Optimal Representation Theorem”) that any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ described above will represent an optimal estimate of $\mathbf{x}^{*}$ in latent state $\mathbf{Z}(t)$. We begin by proving results on optimal estimators in Sections B.1B.2 with identity observation map $f$, developing the linear case of the optimal representation theorem (Theorem B.5) showing that the latent state $\mathbf{Z}(t)$ must encode an estimate of $\mathbf{x}^{*}$ (visualized in Figure S6). We generalize this result any injective observation map $f$ in Section B.3 and derive closed-form solutions for extracting the estimate of $\mathbf{x}^{*}$ from $\mathbf{Z}(t)$. We derive approximation results for $g=\tanh$ in Corollary B.7 and $g=\text{sigmoid}$ in Corollary B.8 that show the representation of $\mathbf{x}^{*}$ in $\mathbf{Z}(t)$ will be linear-affine decodable if $g$ is in the sigmoid family of functions.
@@ B.1 Single Decision Boundary
Note that non-zero noise is required for Lemma B.3 to hold, as zero noise would yield zero probability mass on one side of each decision boundary, meaning that no distance information would be recoverable from $\hat{Y}(t)$ (and Equation 15 would lead to a $0\cdot\infty$ indeterminacy).
@@ B.2 Trilateration via Multiple Decision Boundaries
Let $\hat{\mathbf{Y}}(t)$ denote the vector of classification estimates $\hat{Y}(t)$ from Equation 15. We now have the tools to prove our final result via trilateration. Much like distance information from cell towers can be used to trilaterate [^3] Trilateration differs from triangulation, and it is more frequently used in practice. Triangulation is when one has angle information w.r.t. the cell towers. Usually, this is not available – so one trilaterates their position [57]. This more closely matches our setting, where we just have distances information w.r.t. the decision boundaries and must determine the position. one’s position, we will leverage Lemma B.3 and use distances from decision boundaries $\{(\mathbf{c}_{i},b_{i})\}_{i\in[N_{\textit{task}}]}$ to constrain the positions.
@@
Theorem B.5 boils down to the observation that the confidence associated with each $\hat{Y}_{i}$ in $\hat{\mathbf{Y}}(t)$ are measures of distance between an implied estimate of $\mathbf{x}^{*}$ (denoted $\mu(t)$) and classification boundary $i$ (denoted $(\mathbf{c}_{i},b)$). $\hat{\mathbf{Y}}$ specifies the position of $\hat{\mathbf{X}}=\mu$ via “coordinates” defined by decision boundary normal vectors $\mathbf{c}_{1},\dots,\mathbf{c}_{N_{\textit{task}}}$.
@@ B.3 Optimal Representation Theorem (General Case)
Let $\mathbf{x}^{*}\in\mathbb{R}^{D}$ be a latent representation for linear binary classification task $\mathbf{y}(\mathbf{x}^{*})\in\{0,1\}^{N_{\textit{task}}}$ and $\mathbf{X}(t)=f(\mathbf{x}^{*}+\sigma\mathcal{N}(\mathbf{0},\mathbf{I}_{D}))$ be noisy observations transformed by an injective observation map $f$.
If $\mathbf{C}\in{N_{\textit{task}}\times D}$ is a full-rank matrix representing the decision boundary normal vectors in $\mathbb{R}^{D}$ and $N_{\textit{task}}\geq D$, then any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode an optimal representation of the latent variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$. Furthermore, the mutual information between $\mathbf{x}^{*}$ and $\mathbf{Z}(t)$ is at least as great as the entropy of the theoretically optimal estimator $\mu(t)$, ensuring that all the information of $\mu(t)$ is contained in $\mathbf{Z}(t)$. Consequently, $\mu(t)$ – the optimal estimate of $\mathbf{x}^{*}$ based on $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$ – can be written as a deterministic function (Equation 20) of latent state $\mathbf{Z}(t)$.
@@
Motivated by the similarity between $\Phi(z)$ and sigmoid-like activation functions $g(z)$, we show that the two can be approximately canceled in Equation 20, implying that $\mu(t)$ can be reconstructed with high accuracy with a linear-affine transformation (e.g., linear decoding) when $g=\tanh$ or $g=\text{sigmoid}$. This implies that $\mathbf{Z}(t)$ contains an abstract representation of $\mu(t)$ [59].
@@
Page [60] proposed a simple approximation of $\Phi$ via $\tanh$. Eliminating higher order terms, their approximation is $\Phi(x)\approx\frac{1}{2}\tanh(\sqrt{\frac{2}{\pi}}x)+\frac{1}{2}$. We found the following approximation yielded a superior mean squared error:
@@
Therefore, we have shown that $\mu(t)$ can be expressed as a linear transformation of $\mathbf{Z}(t)$ when the readout function $g$ is sigmoid-like. This confirms the corollary.
Corollary B.8.
For $g=\textup{sigmoid}$, linear scaling by $a_{\sigma}=0.5886$ yields a mean absolute error of 0.0038699 from $Z(t)$ in the range $[-10,10]$, enabling the following accurate linear-affine approximation of $\mu(t)$ from $\mathbf{Z}(t)$ given $g=\textup{sigmoid}$:
$\mu(t)\approx\frac{a_{\sigma}\,\sigma}{\sqrt{t}}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{C}^{\top}\mathbf{Z}(t)+(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{b}$ (25)
Proof.
We found $a_{\sigma}=0.5886$ by computationally minimizing the mean squared error between $\Phi(a_{\sigma}z)$ and $\sigma(z)$ (see sigmoid_approx_gaussianCDF.m in supporting code). Upon computing $a_{\sigma}$ on successively larger optimization bounds $T=1,10,100,...$, we found that $a_{\sigma}$ converged to $0.5886$. We note that sigmoid approximations to the Gaussian distribution CDF have existed in the literature for some time [74]. ∎
Figure S7: Sigmoid ($\sigma(\cdot)$) and $\tanh$ approximations of the normal distribution CDF $\Phi$ via horizontal scaling.
Corollary B.7 and B.8 are visualized in Figure S7, showing the close approximations to the Gaussian CDF.
Corollary B.9.
$N_{\textit{task}}\gg D$ implies orthogonal representations in latent $Z(t)$.
Proof.
Recall equation (3) from the disentangled representation theorem:
$\mu(t)=(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}~\begin{pmatrix}\frac{\sigma}{\sqrt{t}}~\Phi^{-1}\big{(}g(\mathbf{Z}(t))\big{)}+\mathbf{b}\end{pmatrix}$
For sigmoid-like $g$, we can approximate
$\mu(t)\approx\frac{a_{g}\,\sigma}{\sqrt{t}}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{Z}(t)+(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{b}$
The orthogonality of the representations in $\mathbf{Z}(t)$ is therefore governed by the orthogonality of the rows in the matrix $\mathbf{A}:=(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\in\mathbb{R}^{D\times N_{\textit{task}}}$. If $\mathbf{A}\mathbf{A}^{\top}\in\mathbb{R}^{D\times D}$ is diagonal, then the rows of $\mathbf{A}$ are orthogonal.
$\mathbf{A}\mathbf{A}^{\top}=\big{(}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\big{)}\big{(}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\big{)}^{\top}$
$=\big{(}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\big{)}\big{(}\mathbf{C}((\mathbf{C}^{\top}\mathbf{C})^{-1})^{\top}\big{)}$
$=(\mathbf{C}^{\top}\mathbf{C})^{-1}(\mathbf{C}^{\top}\mathbf{C})((\mathbf{C}^{\top}\mathbf{C})^{-1})^{\top}$
$=((\mathbf{C}^{\top}\mathbf{C})^{-1})^{\top}$
$B^{\top}B$ is a symmetric for any matrix B. Recall that the inverse of a symmetric matrix is also symmetric. So $(\mathbf{C}^{\top}\mathbf{C})^{-1}$ is also symmetric. Therefore
$\mathbf{A}\mathbf{A}^{\top}=(\mathbf{C}^{\top}\mathbf{C})^{-1}$
As the columns of $\mathbf{C}$ are high-dimensional randomly sampled vectors, their probability of being orthogonal vanishes as the dimensionality $N_{\textit{task}}$ increases. We can also state the condition in terms of the singular value decomposition (SVD) of $\mathbf{C}=\mathbf{U}\Sigma\mathbf{V}^{T}$ where $\mathbf{U}\in\mathbb{R}^{N_{\textit{task}}\times D}$, $\Sigma=\text{diag}(\sigma_{1},...,\sigma_{D})$, $\mathbf{V}\in\mathbb{R}^{D\times D}$, and $\sigma_{i}$ is the $i$th singular value of $\mathbf{C}$ and $\mathbf{U},\mathbf{V}$ are orthonormal. Then
$\mathbf{A}\mathbf{A}^{\top}=\mathbf{V}\Sigma^{-2}\mathbf{V}^{T}$
If the singular values are approximately uniform $\sigma_{1}\dots\sigma_{D}\approx\sigma$ then
$\mathbf{A}\mathbf{A}^{\top}\approx\mathbf{V}(\frac{1}{\sigma^{2}}\mathbf{I}_{D})\mathbf{V}^{T}$
$\mathbf{A}\mathbf{A}^{\top}\approx\frac{1}{\sigma^{2}}\mathbf{V}\mathbf{V}^{T}=\frac{1}{\sigma^{2}}\mathbf{I}_{D}$
Therefore, uniform singular values in $\mathbf{C}$ is a sufficient condition to guarantee an orthogonal, disentangled representation of $\mu(t)$ in $\mathbf{Z}(t)$ [^4] This uniformity condition in the singular value decomposition is analogous to the outcome of the LM damping technique [45, 50], used for least squares inversion problems in various applications.. ∎
B.4 Suitable Noise Distributions
While the original proof leverages Gaussian noise due to its mathematical convenience, the key property required for the proof is more general. Specifically, the essential requirement is that the marginal posterior distributions along the decision boundary normals $\mathbf{c}_{i}$ have invertible cumulative distribution functions (CDFs), allowing us to recover the distances from the observed classification probabilities.
We will now provide a precise mathematical description of the class of noise distributions where this key property holds, generalizing the disentangled representation theorem beyond Gaussian noise.
Key Noise Property Required for Proof:
Invertibility of the Marginal Posterior CDFs Along Decision Boundary Normals. For each decision boundary normal vector $\mathbf{c}_{i}$, the marginal posterior distribution of $\mathbf{x}^{*}$ projected onto $\mathbf{c}_{i}$ must have an invertible CDF. This allows us to map the observed classification probabilities to unique distances between the estimated mean $\mu(t)$ and the decision boundaries.
Mathematical Description of Suitable Noise Distributions
: Let us define a class of noise distributions $\epsilon(t)$ where the key property holds and is straight forward to solve analytically.
Definition: The proof is immediately generalizable to noise distribution $\epsilon(t)$ if it satisfies the following conditions:
  1. 1.
    Additive noise model:
    $\mathbf{X}(t)=\mathbf{x}^{*}+\epsilon(t)$
    where $\epsilon(t)$ are i.i.d. random vectors.
  2. 2.
    Posterior Distribution Tractability: The posterior distribution $P(\mathbf{x}^{*}|\{\mathbf{X}(s)\}_{s=1}^{t}$ must be analytically tractable or well-approximated, allowing us to compute the posterior mean or maximum a posteriori estimate $\mu(t)$ of $\mathbf{x}^{*}$ and understand its properties.
  3. 3.
    Existence of Invertible Marginal Posterior CDFs: For each decision boundary normal vector $\mathbf{c}_{i}$, the marginal posterior distribution of $\mathbf{c}_{i}^{\top}\mathbf{x}^{*}$ has a continuous and strictly increasing CDF $F_{i}(k)$, which is invertible.
  4. 4.
    Support over $\mathcal{X}$: Let $\mathcal{X}\subseteq\mathbb{R}^{D}$ be the connected subset of allowable $\mathbf{x}^{*}$ values. The noise distribution must have full support over $\mathcal{X}$, ensuring that any real-valued $\mathbf{x}^{*}$ is possible to trilaterate. For our proof, we assume support over the maximally permissible $\mathbb{R}^{D}$ is used.
Implications:
Any suitable noise distribution allows classification task probability $\hat{Y}_{i}(t)$ to be expressed as
$\hat{Y}_{i}(t)=\Pr\{\mathbf{c}_{i}^{\top}\mathbf{x}^{*}>b_{i}|[\mathbf{X}(s)]_{s=1}^{T}\}$
$=1-F_{i}(b_{i}-\mathbf{c}_{i}^{\top}\mu(t))$
where $F_{i}$ is the marginal distribution of $\mathbf{c}_{i}^{\top}\mathbf{x}^{*}$.
Since $F_{i}$ is invertible, we can solve for distance $k_{i}=\mathbf{c}_{i}^{\top}\mu(t)-b_{i}$ as
$k_{i}=F_{i}^{-1}(1-\hat{Y}_{i}(t)).$
This equation allows us to reconstruct decision boundary distances $k_{i}$ from optimal classification probabilities $\hat{Y}_{i}(t)$. Thus the proof via trilateration for the disentangled representation theorem is feasible for any suitable noise distribution $\epsilon(t)$ as described above.
B.5 Examples of Suitable Noise Distributions
B.5.1 Elliptical Distributions
Definition:
A multivariate distribution ([22]) is elliptical if its density function $f(\mathbf{x})$ can be expressed as:
$f(\mathbf{x})=|\boldsymbol{\Sigma}|^{-1/2}g\left((\mathbf{x}-\boldsymbol{\mu})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})\right)$
where $g:[0,\infty)\to[0,\infty)$ is a non-negative function, $\mu\in\mathbb{R}^{D}$ is the location parameter, and $\Sigma\in\mathbb{R}^{D\times D}$ is the scale matrix.
Properties:
  • Symmetric and unimodal around $\boldsymbol{\mu}$.
  • Projections onto any direction $\mathbf{c}_{i}$ yield univariate elliptical distributions.
  • Marginal distributions along $\mathbf{c}_{i}$ have invertible CDFs if $g$ leads to such marginals.
Examples of Suitable Elliptical Distributions:
  • Multi-variate Gaussians with Full-Rank Covariance Matrix (Lemma B.10).
  • Multi-variate T-distributions with Full-Rank Scale Matrix: Heavy-tailed alternative to the Gaussian.
  • Multivariate Laplace Distribution With Full-Rank Scale Matrix: Has exponential tails.
B.5.2 Exponential Power Distributions
Definition: Also known as the generalized Gaussian distribution, defined by the density:
$f(\mathbf{x})=\frac{\beta}{2\alpha\Gamma(1/\beta)}\exp\left(-\left(\frac{|\mathbf{x}-\boldsymbol{\mu}|}{\alpha}\right)^{\beta}\right),$
where $\beta>0$ controls the kurtosis ([7]). Properties:
  • For $\beta=2$, it reduces to the Gaussian distribution.
  • For $\beta=1$, it becomes the Laplace distribution.
  • Symmetric and unimodal.
  • Marginal distributions are of the same form and have invertible CDFs.
Generalizing the Proof
Given the above, the proof can be generalized to any noise distribution $\epsilon(t)$ satisfying the conditions stated. The key steps are as follows:
  1. 1.
    Compute the Posterior Distribution:
    • Since $\epsilon(t)$ is i.i.d., the likelihood function is:
      $P(\{\mathbf{X}(s)\}_{s=1}^{t}\mid\mathbf{x}^{*})=\prod_{s=1}^{t}f_{\epsilon}(\mathbf{X}(s)-\mathbf{x}^{*}).$
    • Without a prior (uniform prior), the posterior is proportional to the likelihood.
    • The posterior distribution $P(\mathbf{x}^{*}\mid\{\mathbf{X}(s)\}_{s=1}^{t})$ can be found (or approximated) based on the noise distribution.
  2. 2.
    Marginalize Along Decision Boundary Normals:
    • For each $\mathbf{c}_{i}$, compute the marginal posterior distribution of $\mathbf{c}_{i}^{\top}\mathbf{x}^{*}$.
    • Due to the symmetry and unimodality of the noise distribution, this marginal will also be symmetric and unimodal.
  3. 3.
    Compute Classification Probabilities: The classification probability is:
    $\hat{Y}_{i}(t)=\Pr\{\mathbf{c}_{i}^{\top}\mathbf{x}^{*}>b_{i}\mid\{\mathbf{X}(s)\}_{s=1}^{t}\}=1-F_{i}(b_{i}-\mathbf{c}_{i}^{\top}\mu(t)),$
    where $F_{i}$ is the marginal posterior CDF of $\mathbf{c}_{i}^{\top}\mathbf{x}^{*}$.
  4. 4.
    Invert Marginal CDFs to Find Distances: Since $F_{i}$ is invertible, we can solve for $k_{i}$:
    $k_{i}=F_{i}^{-1}(1-\hat{Y}_{i}(t)).$
  5. 5.
    Set Up Linear System to Recover $\mu(t)$:
    • The distances $k_{i}$ relate to $\mu(t)$ via:
      $\mathbf{c}_{i}^{\top}\mu(t)=k_{i}+b_{i}.$
    • Collecting all $N_{\text{task}}$ equations:
      $\mathbf{C}\mu(t)=\mathbf{k}+\mathbf{b}.$
  6. 6.
    Solve for $\mu(t)$: If $\mathbf{C}$ is full rank, we can solve for $\mu(t)$:
    $\mu(t)=(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}(\mathbf{k}+\mathbf{b}).$
Implications
  • Optimal Estimator Must Encode $\mu(t)$:
  • The latent state $\mathbf{Z}(t)$ must contain sufficient information to recover $\mu(t)$, as it is essential for optimal classification across all tasks.
  • The theorem holds for any noise distribution satisfying the stated conditions, not just Gaussian noise.
B.5.3 Example with Multivariate t-Distribution
Suppose $\epsilon(t)$ follows a multivariate t-distribution ([39]) with degrees of freedom $\nu>2$:
  1. 1.
    Posterior Distribution: The posterior $P(\mathbf{x}^{*}\mid\{\mathbf{X}(s)\}_{s=1}^{t})$ is also a multivariate t-distribution.
  2. 2.
    Marginal Posterior Distributions: Projections onto $\mathbf{c}_{i}$ yield univariate t-distributions.
  3. 3.
    Invertible Marginal CDFs: The CDF of the t-distribution is known and invertible.
  4. 4.
    Recover Distances: Use the inverse t-CDF to find $k_{i}$:
    $k_{i}=s_{t}\cdot T_{\nu}^{-1}(1-\hat{Y}_{i}(t)),$
    where $s_{t}$ is the scale parameter, and $T_{\nu}^{-1}$ is the inverse CDF of the t-distribution with $\nu$ degrees of freedom.
  5. 5.
    Proceed with the proof: Follow the same steps as before to reconstruct $\mu(t)$.
B.5.4 Example with Anisotropic Gaussian Noise
Lemma B.10.
Suppose $\epsilon(t)$ follows an anisotropic multi-variate Gaussian distribution with full-rank covariance matrix $\Sigma$ and zero mean. Then we can update Equation 14 from Lemma B.2 as
$\hat{Y}(t)$ $\triangleq\Pr(\mathbf{c}^{\top}\mathbf{x}^{*}>b)$ (26)
$=\Phi\begin{pmatrix}\frac{k\sqrt{t}}{\sqrt{\mathbf{c}^{\top}\Sigma\mathbf{c}}}\end{pmatrix}$ (27)
Proof.
Anisotropic noise results in the quadratic form $\mathbf{c}^{\top}\Sigma\mathbf{c}$ in the denominator representing the variance of the marginalized anisotropic noise distribution along decision boundary normal vector $\mathbf{c}_{i}$. As long as $\Sigma$ is non-singular, the remainder of the disentangled representation proof may proceed substituting Equation 14 with Equation 27. ∎
B.5.5 Conclusion
The key property enabling us to recover distances from classification probabilities is the invertibility of the marginal posterior CDFs along the decision boundary normals. This property is not exclusive to Gaussian noise but is shared by a broader class of noise distributions, including but not limited to:
  • Elliptical distributions (e.g., multivariate t-distributions, Laplace distributions).
  • Exponential power distributions.
  • Other symmetric and unimodal distributions with invertible marginals.
Therefore, the proof of the Disentangled Representation Theorem generalizes to any noise distribution satisfying the conditions outlined above. The essential requirement is that we can uniquely map the observed classification probabilities to distances along the normals, allowing us to reconstruct the posterior mean $\mu(t)$ and establish that any optimal estimator must encode this information in its latent state $\mathbf{Z}(t)$.
Correspondence in the structure of noise distribution CDF along marginals $F_{i}$ and point-wise activation functions $g$ (the activation function $\hat{\mathbf{Y}}(t)=g(\mathbf{Z}(t))$).
B.6 Discussion